OpenStudy (anonymous):

Antiderivative of:

OpenStudy (anonymous):

\[\sin ^{3}xcos ^{2}x\]

OpenStudy (anonymous):

\[\Large \int \sin^3 x \cos^2 x dx\] \[\Large \int (\sin ^2 x)\cos^2 x \sin x dx\] \[\Large \int (1 - \cos^2 x)\cos^2 x \sin x dx\]

sam (.sam.):

by parts?

OpenStudy (anonymous):

that should help

OpenStudy (mimi_x3):

u-substitution..

sam (.sam.):

u=cosx du=-sinx dx

sam (.sam.):

\[-\int\limits u^2 \left(1-u^2\right) \, du\] \[\frac{u^5}{5}-\frac{u^3}{3}+c\]